The manager of Beta Company is considering to sell its new product at the price of $500 and the price elasticity of demand at the price range is -0.8. (i) What is the marginal revenue from the sales of the product at the demand point? and (ii) As a consultant to this company, are you going to recommend to the company a higher price or a lower price than $500?

Answers:

a) MR = +$125, and recommend a lower price.

b) MR = -$125 and recommend a higher price.

c) MR = -100, and recommend a higher price.

d) MR = +100, and recommend a lower price.
Answer #1

b) MR = -$125 and recommend a higher price.

(i) P = $500; absolute value of elasticity of demand, e = 0.8
So, MR = P[1-(1/e)] = 500[1-(1/0.8)] = 500(1-1.25) = 500(-0.25) = -125

(ii) As demand is inelastic (because 0.8 < 1) so revenue can be increased by increasing the price. Thus, as a consultant higher price than $500 should be advised.